# Private Online Community Detection for Censored Block Models

### v1

@@ Appendix B Ancillary Lemmas
Let $\{X_{i}\}_{i=1}^{m}$ be a collection of independent random variables, where $p=a\frac{\log n}{n}$, $m=n+o(n)$ and each $X_{i}\overset{i.i.d.}{\sim}p(1-\zeta)\delta_{+1}+p\zeta\delta_{-1}+(1-p)\delta_{0}$. Then,
$\operatorname{Pr}\bigg{(}\sum_{i=1}^{m}X_{i}\leq k_{n}\bigg{)}\leq 1-n^{-a(\sqrt{1-\zeta}-\sqrt{\zeta})^{2}+o(1)},$ (26)
@@
Lemma B.2.
(The Chernoff-Hoeffding Inequality [24]) For a binomial random variable $X$ with $n$ trials and mean $\mu_{X}$, the probability that $X$ deviates from the mean by more than $\beta n$ can be bounded as,
$\operatorname{Pr}(|X-\mu_{X}|\geq\beta n)\leq 2e^{-2\beta^{2}n}\triangleq\delta^{\prime},$ (27)
for any $\beta>0$, and any $\delta^{\prime}\in[0,1)$.
Lemma B.3.
(Bernstein’s Inequality [24]) Let $\{X_{i}\}_{i=1}^{n}$ be a collection of zero-mean independent random variables, where each $X_{i}$ is bounded by $M$, almost surely. Then, for any $t\geq 0$, we have
$\operatorname{Pr}\bigg{(}\sum_{i=1}^{n}X_{i}>t\bigg{)}\leq\exp\left[-\frac{t^{2}}{\sum_{i=1}^{n}E(X_{i}^{2})+\frac{Mt}{3}}\right].$ (28)
@@
(Matrix concentration inequality for dense regime [13]) For a matrix $\mathbf{A}\sim{\mathsf{CBM}}(\bm{\sigma},p,\zeta)$ and $p=\Omega(\log(n)/n)$, we have
$\|\mathbf{A}-\mathds{E}(\mathbf{A})\|^{2}\leq c^{\prime}\log(n)$ (29)
w.h.p. for some constant $c^{\prime}$.
Appendix C Brief Refresher on Differential Privacy
Definition 4 ($(\varepsilon,\delta)$-DP).
A randomized algorithm $\mathcal{A}:\mathcal{D}\rightarrow\mathbb{R}^{d}$ satisfies $(\varepsilon,\delta)$-DP if for any two neighboring datasets $\mathcal{D},\mathcal{D}^{\prime}$ that differ in one record, and any measurable subset $\mathcal{S}\subseteq\text{Range}(\mathcal{A})$, we have
$\operatorname{Pr}(\mathcal{A}(\mathcal{D})\in\mathcal{S})\leq e^{\varepsilon}\operatorname{Pr}(\mathcal{A}(\mathcal{D}^{\prime})\in\mathcal{S})+\delta,$ (30)
where $\epsilon$ is coined as privacy budget and $\delta$ represents failure probability. It worth noting that smaller $\epsilon$, $\delta$ correspond to stronger privacy guarantee and vice versa. The setting when $\delta=0$ is referred as pure $\varepsilon$-DP.
Definition 5.
(Sequential Composition of DP [12]) If a randomized mechanism $\mathcal{M}_{i}$ satisfies $(\epsilon_{i},\delta_{i})$-DP for $i=1,2,\cdots,n$, then for any $\delta_{i}\in(0,1]$, their composition $\mathcal{M}$ defined by $\mathcal{M}\triangleq(\mathcal{M}_{1},\mathcal{M}_{2},\cdots,\mathcal{M}_{T})$ satisfies $(\sum_{i=1}^{T}\epsilon_{i},\sum_{i=1}^{T}\delta_{i})$-DP.
Definition 6.
(Advanced Composition of DP [12]) If a sequence of randomized mechanisms $\mathcal{M}_{i}$ each satisfies $(\epsilon,\delta)$-differential privacy for $i=1,2,\cdots,T$, then their composition $\mathcal{M}$, defined by $\mathcal{M}\triangleq(\mathcal{M}_{1},\mathcal{M}_{2},\cdots,\mathcal{M}_{T})$, satisfies $\left(\sqrt{2T\ln(1/\delta^{\prime})}\epsilon+T\epsilon(e^{\epsilon}-1),T\delta+\delta^{\prime}\right)$-differential privacy for any $\delta^{\prime}>0$.
Appendix D Deriving Recovery Condition for Exact Recovery
In this subsection, we analyze the error probability of the semidefinite relaxation of the ML estimator as shown in eqn. (6). Moreover, we derive the necessary conditions for the optimal solution. The Lagrangian function is written as follows:
$\mathcal{L}({\mathbf{A}},\mathbf{Y},\mathbf{S},\mathbf{D})=\operatorname{tr}({\mathbf{A}}\mathbf{Y})+\operatorname{tr}(\mathbf{S}\mathbf{Y})-\operatorname{tr}(\mathbf{D}(\mathbf{Y}-\mathbf{I})),$ (31)
@@ Appendix D Deriving Recovery Condition for Exact Recovery
$\nabla_{\mathbf{Y}}\mathcal{L}={\mathbf{A}}+\mathbf{S}-\mathbf{D}=\mathbf{0}.$ (32)
In order to satisfy the first order stationary condition, we must satisfy the following (deterministic) condition:
$\mathbf{S}^{*}=\mathbf{D}^{*}-{\mathbf{A}},$ (33)
From the KKT conditions, we have the following:
$\mathbf{D}^{*}\mathbf{Y}^{*}=\mathbf{D}^{*}\mathbf{I}\Rightarrow\operatorname{tr}(\mathbf{D}^{*}(\mathbf{Y}^{*}-\mathbf{I}))=0,$ (34)
@@
where in eqn. (35), ${\bm{\sigma}}^{*}$ is the null space of $\mathbf{S}^{*}$. In order to ensure that $\mathbf{Y}^{*}$ is the unique solution, we require that $\lambda_{2}(\mathbf{S}^{*})>0$, i.e., the second smallest eigenvalue of $\mathbf{S}$. This comes from the rank-nullity theorem, i.e.,
$\text{rank}(\mathbf{S}^{*})+\text{null}(\mathbf{S}^{*})=n\Rightarrow\text{rank}(\mathbf{S}^{*})=n-1.$ (36)
To this end, we have the following:
@@
$=\operatorname{tr}((\mathbf{S}^{*}+{\mathbf{A}})\mathbf{Y}^{*})=\operatorname{tr}({\mathbf{A}}\mathbf{Y}^{*}).$ (37)
@@
$\operatorname{Pr}\left[\inf_{\mathbf{x}:\|\mathbf{x}\|=1,\mathbf{x}\perp\bm{\sigma}^{*}}\mathbf{x}^{T}\mathbf{S}^{*}\mathbf{x}>0\right]\geq 1-o(1).$ (38)
@@
$\mathds{E}[{\mathbf{A}}]$ $=(1-2\zeta)p(\mathbf{Y}^{*}-\mathbf{I}).$ (39)
Now, for any $\mathbf{x}$ such that $\|\mathbf{x}\|=1,\mathbf{x}\perp\bm{\sigma}^{*}$ (i.e., $\mathbf{x}^{T}\bm{\sigma}^{*}$ = 0), we proceed with the following set of steps:
@@
=$\overset{(a)}{=}\mathbf{x}^{T}\mathbf{D}^{*}\mathbf{x}+(1-2\zeta)p-\|{\mathbf{A}}-\mathds{E}[{\mathbf{A}}]\|$
@@
where step (a) holds from the unique optimality condition in (35). Note that $d_{i}$ is equal in distribution to $\sum_{i=1}^{n-1}X_{i}$, where $X_{i}\overset{i.i.d.}{\sim}p(1-\zeta)\delta_{+1}+p\zeta\delta_{-1}+(1-p)\delta_{0}$. Step (b) follows that $\|\tilde{\mathbf{A}}-\mathds{E}[\tilde{\mathbf{A}}]\|\leq\tilde{c}^{\prime}\sqrt{\log(n)}$ with high probability, as convincingly demonstrated in [17]. The invocation of Lemma B.1 and the subsequent application of the union bound unequivocally validate the condition.
D.1 The Synergy of Combining Multiple Adjacency Matrices
@@
$\hat{\mathbf{Y}}=\max_{\mathbf{Y}}$ $\ \text{tr}\left(\sum_{t=1}^{w}{\mathbf{A}}_{t}\mathbf{Y}\right)$
s.t. $\ \mathbf{Y}\succcurlyeq\mathbf{0},Y_{i,i}=1,\forall i\in[n].$
The Lagrangian function is written as follows:
$\mathcal{L}({\mathbf{A}}_{1},\ldots,{\mathbf{A}}_{w},\mathbf{Y},\mathbf{S},\mathbf{D})=\operatorname{tr}\left(\sum_{t=1}^{w}{\mathbf{A}}_{t}\mathbf{Y}\right)+\operatorname{tr}(\mathbf{S}\mathbf{Y})-\operatorname{tr}(\mathbf{D}(\mathbf{Y}-\mathbf{I})),$ (41)
@@
$\nabla_{\mathbf{Y}}\mathcal{L}=\sum_{t=1}^{w}{\mathbf{A}}_{t}+\mathbf{S}-\mathbf{D}=\mathbf{0}.$ (42)
In order to satisfy the first order stationary condition, we must satisfy the following (deterministic) condition:
$\mathbf{S}^{*}=\mathbf{D}^{*}-\sum_{t=1}^{w}{\mathbf{A}}_{t},$ (43)
Using Lemma 9 in [17], we have when ${\mathbf{S}}^{\ast}={\mathbf{D}}^{\ast}-\sum_{t=1}^{w}{\mathbf{A}}_{t}$ satisfies ${\mathbf{S}}^{\ast}\succcurlyeq\mathbf{0}$, $\lambda_{2}({\mathbf{S}}^{\ast})>0$ (second smallest eigenvalue) and
${\mathbf{S}}^{\ast}\bm{\sigma}^{\ast}=0,$
then $\hat{\mathbf{Y}}=\mathbf{Y}^{\ast}=\bm{\sigma}^{\ast}(\bm{\sigma}^{\ast})^{\top}$ is the unique solution. Therefore, it suffices to show ${\mathbf{S}}^{\ast}\succcurlyeq\mathbf{0}$ and $\lambda_{2}({\mathbf{S}}^{\ast})>0$. We let
${\mathbf{D}}^{\ast}=\text{diag}\{d_{i}^{\ast}\},\quad d_{i}^{\ast}=\sum_{t=1}^{w}\sum_{j=1}^{n}{\mathbf{A}}_{t}(i,j)\sigma_{i}^{\ast}\sigma_{j}^{\ast}.$
By definition we have $d_{i}^{\ast}\sigma_{i}^{\ast}=\sum_{t=1}^{w}\sum_{j=1}^{n}{\mathbf{A}}_{t}(i,j)\sigma_{j}^{\ast}$, i.e., ${\mathbf{D}}^{\ast}\bm{\sigma}^{\ast}=(\sum_{t=1}^{w}{\mathbf{A}}_{t})\bm{\sigma}^{\ast}$, thus ${\mathbf{S}}^{\ast}\bm{\sigma}^{\ast}=0$. It remains to show that
$\operatorname{Pr}\left[\inf_{\mathbf{x}:\|\mathbf{x}\|=1,\mathbf{x}\perp\bm{\sigma}^{*}}\mathbf{x}^{T}\mathbf{S}^{*}\mathbf{x}>0\right]\geq 1-o(1).$ (44)
@@
$\mathds{E}[\sum_{t=1}^{w}{\mathbf{A}}_{t}]$ $=w(1-2\zeta)p(\mathbf{Y}^{*}-\mathbf{I}).$ (45)
Now, for any $\mathbf{x}$ such that $\|\mathbf{x}\|=1,\mathbf{x}\perp\bm{\sigma}^{*}$ (i.e., $\mathbf{x}^{T}\bm{\sigma}^{*}$ = 0), we proceed with the following set of steps:
$\mathbf{x}^{T}\mathbf{S}^{*}\mathbf{x}$ $=\mathbf{x}^{T}\mathbf{D}^{*}\mathbf{x}-\mathbf{x}^{T}\mathds{E}[\sum_{t=1}^{w}{\mathbf{A}}_{t}]\mathbf{x}-\mathbf{x}^{T}(\sum_{t=1}^{w}{\mathbf{A}}_{t}-\mathds{E}[\sum_{t=1}^{w}{\mathbf{A}}_{t}])\mathbf{x}$
$\geq\mathbf{x}^{T}\mathbf{D}^{*}\mathbf{x}-w(1-2\zeta)p\mathbf{x}^{T}\mathbf{Y}^{*}\mathbf{x}+w(1-2\zeta)p-\|\sum_{t=1}^{w}{\mathbf{A}}_{t}-\mathds{E}[\sum_{t=1}^{w}{\mathbf{A}}_{t}]\|$
=$\overset{(a)}{=}\mathbf{x}^{T}\mathbf{D}^{*}\mathbf{x}+w(1-2\zeta)p-\|\sum_{t=1}^{w}{\mathbf{A}}_{t}-\mathds{E}[\sum_{t=1}^{w}{\mathbf{A}}_{t}]\|$
$\geq\min_{i\in[n]}d_{i}^{*}+w(1-2\zeta)p-\|\sum_{t=1}^{w}{\mathbf{A}}_{t}-\mathds{E}[\sum_{t=1}^{w}{\mathbf{A}}_{t}]\|,$ (46)
where step (a) holds since $\mathbf{x}\perp\bm{\sigma}^{*}$. Note that each $d_{i}^{\ast}$ is equal in distribution to $\sum_{j=1}^{w(n-1)}X_{j}$, where $X_{j}\overset{i.i.d.}{\sim}p(1-\zeta)\delta_{+1}+p\zeta\delta_{-1}+(1-p)\delta_{0}$.
First, we have for $m=w(n-1)$, and $k_{n}>0$,

### v2

@@ B-A Ancillary Lemmas
Let $\{X_{i}\}_{i=1}^{m}$ be a collection of independent random variables, where $p=a\frac{\log n}{n}$, $m=\rho n+o(n)$ for some $\rho>0$ as $n\to\infty$, and each $X_{i}\overset{i.i.d.}{\sim}p(1-\zeta)\delta_{+1}+p\zeta\delta_{-1}+(1-p)\delta_{0}$. Then,
$\operatorname{Pr}\bigg(\sum_{i=1}^{m}X_{i}\leq k_{n}\bigg)\leq n^{-\rho a(\sqrt{1-\zeta}-\sqrt{\zeta})^{2}+o(1)},$
@@
(Matrix concentration inequality for dense regime [20]) For a matrix $\mathbf{A}\sim{\mathsf{CBM}}(\bm{\sigma},p,\zeta)$ and $p=\Omega(\log(n)/n)$, we have for any $c>0$, there exists $c^{\prime}>0$ such that
$\operatorname{Pr}\bigg(\|\mathbf{A}-\mathds{E}(\mathbf{A})\|\leq c^{\prime}\sqrt{np}\bigg)\geq 1-n^{-c}.$
B-B Deriving Recovery Condition for Exact Recovery
In this subsection, we review the derivation of necessary conditions, $a(\sqrt{1-\zeta}-\sqrt{\zeta})^{2}>1$, for exact community recovery using SDP estimate (3), under non-private scenarios, as shown in [20]. This derivation forms the foundational basis for our subsequent proofs. We also shed light on the necessary condition when multiple graphs are used.
First of all, for the semidefinite program in (3), the Lagrangian function is written as follows,
$\mathcal{L}({\mathbf{A}},\mathbf{Y},\mathbf{S},\mathbf{D})=\operatorname{tr}({\mathbf{A}}\mathbf{Y})+\operatorname{tr}(\mathbf{S}\mathbf{Y})-\operatorname{tr}(\mathbf{D}(\mathbf{Y}-\mathbf{I})),$
@@ B-B Deriving Recovery Condition for Exact Recovery
$\nabla_{\mathbf{Y}}\mathcal{L}={\mathbf{A}}+\mathbf{S}-\mathbf{D}=\mathbf{0}.$
In order to satisfy the first order stationary condition, we must satisfy the (deterministic) condition
$\mathbf{S}^{*}=\mathbf{D}^{*}-{\mathbf{A}},$
From the KKT conditions, we have the following,
$\mathbf{D}^{*}\mathbf{Y}^{*}=\mathbf{D}^{*}\mathbf{I}\Rightarrow\operatorname{tr}(\mathbf{D}^{*}(\mathbf{Y}^{*}-\mathbf{I}))=0,$
@@
where in Eqn. (28), $\bm{\sigma}$$^{*}$ is the null space of $\mathbf{S}^{*}$. In order to ensure that $\mathbf{Y}^{*}$ is the unique solution, we require that $\lambda_{2}(\mathbf{S}^{*})>0$, i.e., the second smallest eigenvalue of $\mathbf{S}$ is positive. This comes from the rank-nullity theorem, i.e.,
$\text{rank}(\mathbf{S}^{*})+\text{null}(\mathbf{S}^{*})=n\Rightarrow\text{rank}(\mathbf{S}^{*})=n-1.$
To this end, we have the following,
@@
$=\operatorname{tr}((\mathbf{S}^{*}+{\mathbf{A}})\mathbf{Y}^{*})=\operatorname{tr}({\mathbf{A}}\mathbf{Y}^{*}).$
@@
$\operatorname{Pr}\left[\inf_{\mathbf{x}:\|\mathbf{x}\|=1,\mathbf{x}\perp\bm{\sigma}^{*}}\mathbf{x}^{T}\mathbf{S}^{*}\mathbf{x}>0\right]\geq 1-o(1).$
@@
$\mathds{E}[{\mathbf{A}}]$ $=(1-2\zeta)p(\mathbf{Y}^{*}-\mathbf{I}).$
Now, for any $\mathbf{x}$ such that $\|\mathbf{x}\|=1,\mathbf{x}\perp\bm{\sigma}^{*}$ (i.e., $\mathbf{x}^{T}\bm{\sigma}^{*}$ = 0), we proceed with the following set of steps,
@@
$\overset{(a)}{=}\mathbf{x}^{T}\mathbf{D}^{*}\mathbf{x}+(1-2\zeta)p-\|{\mathbf{A}}-\mathds{E}[{\mathbf{A}}]\|$
@@
where step (a) holds from the unique optimality condition in (28). Note that $d_{i}$ is equal in distribution to $\sum_{i=1}^{n-1}X_{i}$, where $X_{i}\overset{i.i.d.}{\sim}p(1-\zeta)\delta_{+1}+p\zeta\delta_{-1}+(1-p)\delta_{0}$. Step (b) follows that $\|{\mathbf{A}}-\mathds{E}[{\mathbf{A}}]\|\leq c^{\prime}\sqrt{\log(n)}$ with high probability by Lemma B.2 when $p=a\frac{\log n}{n}$. Using Lemma B.1 and the subsequent application of the union bound yields that with probability at least $1-n^{1-a(\sqrt{1-{\zeta}}-\sqrt{{\zeta}})^{2}+o(1)}$, we have
$\min_{i}d_{i}^{\ast}\geq\frac{\log n}{\log\log n},$
and $\frac{\log n}{\log\log n}+(1-2\zeta)p-c^{\prime}\sqrt{a\log n}>0$ for $n$ sufficiently large. Therefore if $a(\sqrt{1-{\zeta}}-\sqrt{{\zeta}})^{2}>1$, the model satisfies the exact recovery condition.
Remark B.1 (The Synergy of Combining Multiple Adjacency Matrices).
@@
$\hat{\mathbf{Y}}=\max_{\mathbf{Y}}$ $\ \text{tr}\left(\sum_{t=1}^{w}{\mathbf{A}}_{t}\mathbf{Y}\right)\ \text{s.t.}\ \mathbf{Y}\succcurlyeq\mathbf{0},Y_{i,i}=1,\forall i\in[n].$
The Lagrangian function is similarly written as
$\mathcal{L}({\mathbf{A}}_{1},\ldots,{\mathbf{A}}_{w},\mathbf{Y},\mathbf{S},\mathbf{D})=\operatorname{tr}\left(\sum_{t=1}^{w}{\mathbf{A}}_{t}\mathbf{Y}\right)+\operatorname{tr}(\mathbf{S}\mathbf{Y})$
$\hskip 120.0pt-\operatorname{tr}(\mathbf{D}(\mathbf{Y}-\mathbf{I})),$
@@
$\nabla_{\mathbf{Y}}\mathcal{L}=\sum_{t=1}^{w}{\mathbf{A}}_{t}+\mathbf{S}-\mathbf{D}=\mathbf{0}.$
In order to satisfy the first order stationary condition, we must satisfy the following (deterministic) condition
$\mathbf{S}^{*}=\mathbf{D}^{*}-\sum_{t=1}^{w}{\mathbf{A}}_{t}.$
Similar to the single observation case, we have that when ${\mathbf{S}}^{\ast}\succcurlyeq\mathbf{0}$, $\lambda_{2}({\mathbf{S}}^{\ast})>0$, and ${\mathbf{S}}^{\ast}\bm{\sigma}^{\ast}=0$, then $\hat{\mathbf{Y}}=\mathbf{Y}^{\ast}=\bm{\sigma}^{\ast}(\bm{\sigma}^{\ast})^{T}$ is the unique solution. We let
${\mathbf{D}}^{\ast}=\text{diag}\{d_{i}^{\ast}\},\qquad d_{i}^{\ast}=\sum_{t=1}^{w}\sum_{j=1}^{n}{\mathbf{A}}_{t}(i,j)\sigma_{i}^{\ast}\sigma_{j}^{\ast}.$
By definition we have $d_{i}^{\ast}\sigma_{i}^{\ast}=\sum_{t=1}^{w}\sum_{j=1}^{n}{\mathbf{A}}_{t}(i,j)\sigma_{j}^{\ast}$, i.e., ${\mathbf{D}}^{\ast}\bm{\sigma}^{\ast}=(\sum_{t=1}^{w}{\mathbf{A}}_{t})\bm{\sigma}^{\ast}$, thus ${\mathbf{S}}^{\ast}\bm{\sigma}^{\ast}=0$. It remains to show
$\operatorname{Pr}\left[\inf_{\mathbf{x}:\|\mathbf{x}\|=1,\mathbf{x}\perp\bm{\sigma}^{*}}\mathbf{x}^{T}\mathbf{S}^{*}\mathbf{x}>0\right]\geq 1-o(1).$
@@
$\mathds{E}[\sum_{t=1}^{w}{\mathbf{A}}_{t}]$ $=w(1-2\zeta)p(\mathbf{Y}^{*}-\mathbf{I}).$
Now, for any $\mathbf{x}$ such that $\|\mathbf{x}\|=1,\mathbf{x}\perp\bm{\sigma}^{*}$ (i.e., $\mathbf{x}^{T}\bm{\sigma}^{*}$ = 0), we proceed similarly as (29),
$\mathbf{x}^{T}\mathbf{S}^{*}\mathbf{x}=\mathbf{x}^{T}\mathbf{D}^{*}\mathbf{x}-\mathbf{x}^{T}\mathds{E}[\sum_{t=1}^{w}{\mathbf{A}}_{t}]\mathbf{x}$
$\hskip 50.0pt-\mathbf{x}^{T}(\sum_{t=1}^{w}{\mathbf{A}}_{t}-\mathds{E}[\sum_{t=1}^{w}{\mathbf{A}}_{t}])\mathbf{x}$
$\geq\min_{i\in[n]}d_{i}^{*}+w(1-2\zeta)p-\|\sum_{t=1}^{w}{\mathbf{A}}_{t}-\mathds{E}[\sum_{t=1}^{w}{\mathbf{A}}_{t}]\|.$ (30)
Note that each $d_{i}^{\ast}$ is equal in distribution to $\sum_{j=1}^{w(n-1)}X_{j}$, where $X_{j}\overset{i.i.d.}{\sim}p(1-\zeta)\delta_{+1}+p\zeta\delta_{-1}+(1-p)\delta_{0}$. By Lemma B.1 and the union bound, we have with probability at least $1-n^{1-wa(\sqrt{1-{\zeta}}-\sqrt{{\zeta}})^{2}+o(1)}$, $\min_{i}d_{i}^{\ast}\geq\frac{\log n}{\log\log n}$. Second, we have
$\|\sum_{t=1}^{w}{\mathbf{A}}_{t}-\mathds{E}[\sum_{t=1}^{w}{\mathbf{A}}_{t}]\|\leq\sum_{t=1}^{w}\|{\mathbf{A}}_{t}-\mathds{E}[{\mathbf{A}}_{t}]\|,$
by Lemma B.2 we have $\operatorname{Pr}\{\|{\mathbf{A}}_{t}-\mathds{E}[{\mathbf{A}}_{t}]\|\leq c^{\prime}\sqrt{a\log n}\}\geq 1-n^{c}$ for $p=a\frac{\log n}{n}$, thus we have with probability at least $1-wn^{-c}$,
$\|\sum_{t=1}^{w}{\mathbf{A}}_{t}-\mathds{E}[\sum_{t=1}^{w}{\mathbf{A}}_{t}]\|\leq wc^{\prime}\sqrt{a\log n}.$
Combing these together, with probability at least $1-n^{1-wa(\sqrt{1-{\zeta}}-\sqrt{{\zeta}})^{2}+o(1)}-wn^{-c}$, we have (30) $\geq\frac{\log n}{\log\log n}+w(1-2\zeta)p-wc^{\prime}\sqrt{a\log n}>0$ for $n$ sufficiently large and $w=o(\sqrt{\log n})$. And therefore the model satisfies the exact recovery condition if $wa(\sqrt{1-{\zeta}}-\sqrt{{\zeta}})^{2}>1$ and $w=o(\sqrt{\log n})$.
Appendix C Proofs for Section IV-A
C-A Proof of Theorem IV.1
In this subsection, we derive the condition for exact recovery under the privacy mechanism.
Lemma C.1.